Alex has saved $8.80 in in nickels dimes and quarters. If he has 4 nickels fewer than 3 times the number of dimes and one quarter more than 3/4 the number of dimes. How many coins of each type does he have?
Answer by lwsshak3(11628)

Alex has saved $8.80 in in nickels dimes and quarters. If he has 4 nickels fewer than 3 times the number of dimes and one quarter more than 3/4 the number of dimes. How many coins of each type does he have?
***
let x=number of dimes
3x-4=number of nickels
(3/4)x+1=number of quarters
...
.10x+.05(3x-4)+.25(3x/4+1)=8.80
.10x+.15x-.20+.75x/4+.25=8.80
LCD:4
.40x+.60x-.80+.75x+1.00=35.20
1.75x+.20=35.20
1.75x=35.00
x=20
number of dimes=20
number of nickels=56
number of quarters=16